# biblio.bib

### v1

@@ 1 Introduction
In this work, we present Sei Giga, the first multi-proposer EVM layer-1 blockchain. Sei Giga uses Autobahn [giridharan2025autobahnseamlesshighspeed] as a consensus protocol to allow for $5$ gigagas throughput [5giga] and sub $400$ms finality under standard BFT-style security assumptions. Additionally, Sei Giga uses a custom EVM execution client that has been built from scratch, a new storage layer, and asynchronous accumulator-based state commitments. As a result, Sei Giga is able to scale to allow for web2 style usage while ensuring quick finality in settings where it is necessary such as trading, on a verifiable and secure public ledger. Sei Giga is a decentralized, permissionless Proof-of-Stake EVM chain capable of running standard EVM smart contracts written in typical EVM languages such as Solidity and Vyper. Sei Giga’s EVM support is equivalent to mainchain Ethereum with the exception of EIP-4844, PREVRANDAO, the state root, the block gas limit, and the transaction fee mechanism, and Sei Giga will continue to support EVM upgrades to maintain near parity. As is standard in Proof-of-Stake blockchains, Sei Giga is maintained by a set of staked nodes which act as validators and executors in the network responsible for ensuring consensus is met and blocks are run; the nodes are staked and eligible for block rewards to ensure an economic incentive to keep the chain running, and subject to slashing of the staked coins in the case of malicious behavior.
@@ 1.1 Comparison to Sei v2
Autobahn achieves over $50$× higher throughput than Tendermint by fundamentally rethinking how consensus and data availability are managed in a distributed system. In Tendermint, each block is processed sequentially: A single leader proposes a block, validators must download and fully execute the block transactions, and then consensus is reached in multiple rounds before the next block is proposed. This tight coupling between data dissemination, execution, and consensus means that every block decision incurs significant overhead and waiting time, resulting in a throughput limited by the speed of one proposer and the time taken to complete all three rounds of communication.
@@
Overall, by decoupling execution from consensus, enabling multi-proposer parallelism, committing multiple blocks in a single tip cut, reducing the number of communication rounds, and separating data availability from the critical consensus path, the Sei Giga upgrade brings significant improvements over Sei v2. The improved execution client achieves a $40$x performance boost with a wholly rebuilt client from scratch, enabling $200$k TPS compared to the previous $5$k TPS. The faster consensus algorithm reduces voting rounds from $3$ to $1.5$, resulting in a $2$x improvement in consensus efficiency. The novel data availability layer with its multi-producer design delivers a $70$x improvement in block production with $180$ blocks versus $2.5$ previously while still ensuring strong Byzantine Fault Tolerance.
@@ 2.3 Pipelining and encoding
12:         Read each storage key from $payload$
@@ 3.4 Consensus Layer (Cut of Tips)
    Confirm: If the leader gathers only $(n-f)$ votes, it enters a confirm phase and waits for additional acknowledgments until a commit certificate is achieved with $12f+$ 1 confirm messages.
Pipelining Benefit:
Through the use of quadratic communication and pipelining, Sei Giga achieves a slow path round trip of 1.5 rather than 2.5. That is to say that the leader sends a proposal for tip cut $N$, triggering nodes to send their prepare votes. Once enough votes are collected, they form a PrepareQC. After that, nodes send their commit votes, and once enough are collected, a CommitQC is formed, finalizing the blocks in $N$. While $N$ is in its commit phase, the leader for tip cut $N+1$ may start sending its proposal. In other words, even though each block goes through two distinct rounds, pipelining allows the next block’s proposal to start during the commit phase of the previous block, effectively reducing the overall latency to 1.5 rounds.
@@ 6 Economics
Sei Giga has a native coin called SEI with a total supply of $10$ billion, $10\_000\_000\_000$. The Sei coin is used to pay for gas and other associated transaction and usage costs, as well as for staking and validator rewards. Each block receives a block reward in Sei which is not paid to the block producer but the sum of all rewards in an epoch is paid equally to allowed validators and their delegators, $r=\frac{\sum_{b\in e}b_{r}}{|V|}$. Each validator must stake Sei to run a validator node and the stake is subject to a 21 day bonding period, the 21 day period also applies to delegated stake, chosen to allow for slashing of malicious behaviour by nodes. Slashing penalties apply to liveness, and double signing of blocks at the same height. Slashing penalties vary by the severity of the action, such that double signing incurs a greater penalty than liveness, and increase in severity for future penalties.
@@ 6.1 Governance
[…]

### v2

# Benjamin Marsh Steven Landers Jayendra Jog

@@ 1 Introduction
In this work, we present Sei Giga, the first multi-proposer EVM layer-1 blockchain. Sei Giga uses Autobahn [giridharan2025autobahnseamlesshighspeed] as a consensus protocol to allow for $5$ gigagas throughput [5giga] and sub 400400ms finality under standard BFT-style security assumptions. Additionally, Sei Giga uses a custom EVM execution client that has been built from scratch, a new storage layer, and asynchronous accumulator-based state commitments. As a result, Sei Giga is able to scale to allow for web2 style usage while ensuring quick finality in settings where it is necessary such as trading, on a verifiable and secure public ledger. Sei Giga is a decentralized, permissionless Proof-of-Stake EVM chain capable of running standard EVM smart contracts written in typical EVM languages such as Solidity and Vyper. Sei Giga’s EVM support is equivalent to mainchain Ethereum with the exception of EIP-4844, PREVRANDAO, the state root, the block gas limit, and the transaction fee mechanism, and Sei Giga will continue to support EVM upgrades to maintain near parity. As is standard in Proof-of-Stake blockchains, Sei Giga is maintained by a set of staked nodes which act as validators and executors in the network responsible for ensuring consensus is met and blocks are run; the nodes are staked and eligible for block rewards to ensure an economic incentive to keep the chain running, and subject to slashing of the staked coins in the case of malicious behavior.
@@ 1.1 Comparison to Sei v2
Autobahn achieves over 5050× higher throughput than Tendermint by fundamentally rethinking how consensus and data availability are managed in a distributed system. In Tendermint, each block is processed sequentially: A single leader proposes a block, validators must download and fully execute the block transactions, and then consensus is reached in multiple rounds before the next block is proposed. This tight coupling between data dissemination, execution, and consensus means that every block decision incurs significant overhead and waiting time, resulting in a throughput limited by the speed of one proposer and the time taken to complete all three rounds of communication.
@@
Overall, by decoupling execution from consensus, enabling multi-proposer parallelism, committing multiple blocks in a single tip cut, reducing the number of communication rounds, and separating data availability from the critical consensus path, the Sei Giga upgrade brings significant improvements over Sei v2. The improved execution client achieves a 4040x performance boost with a wholly rebuilt client from scratch, enabling 200200k TPS compared to the previous 55k TPS. The faster consensus algorithm reduces voting rounds from $3$ to $1.5$, resulting in a 22x improvement in consensus efficiency. The novel data availability layer with its multi-producer design delivers a 7070x improvement in block production with $180$ blocks versus $2.5$ previously while still ensuring strong Byzantine Fault Tolerance.
@@ 2.3 Pipelining and encoding
4:Read and store $\mathit{Type}$, $\mathit{ChainID}$, $\mathit{Sender}$, $\mathit{To}$, $\mathit{Value}$, and other fields by consuming the appropriate number of bytes from payloadpayload.
@@
12:   Read each storage key from payloadpayload
@@ 3.4 Consensus Layer (Cut of Tips)
    Confirm: If the leader gathers only (nf)(n-f) votes, it enters a confirm phase and waits for additional acknowledgments until a commit certificate is achieved with $2f+1$ confirm messages.
Pipelining Benefit. Through the use of quadratic communication and pipelining, Sei Giga achieves a slow path round trip of 1.5 rather than 2.5. That is to say that the leader sends a proposal for tip cut $N$, triggering nodes to send their prepare votes. Once enough votes are collected, they form a PrepareQC. After that, nodes send their commit votes, and once enough are collected, a CommitQC is formed, finalizing the blocks in $N$. While $N$ is in its commit phase, the leader for tip cut $N+1$ may start sending its proposal. In other words, even though each block goes through two distinct rounds, pipelining allows the next block’s proposal to start during the commit phase of the previous block, effectively reducing the overall latency to 1.5 rounds.
@@ 3.5 System Model, Security, and Reliable Inclusion
3.6 Future consensus improvements
Autobahn-style consensus protocols improve on traditional sequential models by tolerating offline lanes without halting block production or consensus. However, the reliance on a single tip-cut leader makes the system brittle: if the leader goes offline, finality stalls. With $f$ Byzantine nodes and five timeouts at two seconds each, p50 latency degrades to over 11s (compared to 0.25s p50 and 3s p99 in the non-Byzantine case), due to timeouts. This erodes the UX benefits of Giga, since inclusion and execution latency both rise. For some users, the impact is catastrophic. For example, a trader bidding on a cross-chain asset may see their utility fall to zero if consensus latency allows prices on other chains to move beyond their valuation. Thus, while timeouts are infrequent, their user impact is non-negligible. A leaderless or backup-leader variant of Autobahn removes this single point of failure, preserving throughput and low median latency while substantially tightening worst case performance under degraded or adversarial conditions. In real-world networks we have observed a missed slot rate of roughly $\frac{1}{90}$, with some validators missing 15-20% of slots over multi-day periods.
@@ 6 Economics
Sei Giga has a native coin called SEI with a total supply of $10$ billion, 10_000_000_00010\_000\_000\_000. The Sei coin is used to pay for gas and other associated transaction and usage costs, as well as for staking and validator rewards. Each block receives a block reward in Sei which is not paid to the block producer but the sum of all rewards in an epoch is paid equally to allowed validators and their delegators, $r=\frac{\sum_{b\in e}b_{r}}{|V|}$. Each validator must stake Sei to run a validator node and the stake is subject to a 21 day bonding period, the 21 day period also applies to delegated stake, chosen to allow for slashing of malicious behaviour by nodes. Slashing penalties apply to liveness, and double signing of blocks at the same height. Slashing penalties vary by the severity of the action, such that double signing incurs a greater penalty than liveness, and increase in severity for future penalties.
@@ 6.1 Governance
Governance on Sei is a community-driven process, that mirrors the approach taken in v2, that lets stakers propose and vote on network changes. After a proposal is submitted and the required deposit is met to discourage spam, it moves to a voting phase where stakers weigh in. If quorum and majority requirements are satisfied, the proposal passes; otherwise, it fails and deposits may be burned. Passed proposals that change the chain’s parameters or software are automatically enacted, while more general proposals require manual execution. All governance data is publicly accessible on Seiscan or via direct blockchain queries.